OpenStudy (anonymous):

find the value of the expression when y=-3 y^3/y^2-4

OpenStudy (anonymous):

y^3/(y^2 - 4) =-(27/5) when y = -3

OpenStudy (jdoe0001):

\(\bf \cfrac{{\color{red}{ y}}^3}{{\color{red}{ y}}^2-4}\qquad y=-3\implies \cfrac{({\color{red}{ -3}})^3}{({\color{red}{ -3}})^2-4}\)

jimthompson5910 (jim_thompson5910):

Replace ALL copies of 'y' with -3 (since y = -3) and then evaluate. \[\Large \frac{y^3}{y^2-4}\] \[\Large \frac{(-3)^3}{(-3)^2-4}\] \[\Large \frac{-27}{9-4}\] \[\Large -\frac{27}{5}\] So, \[\Large \frac{y^3}{y^2-4} = -\frac{27}{5} \ \text{when } y = -3\]